f(x) = 2x-1
find f(-9)

Answer by bucky(2189)

You are given the expression:
f(x) = 2x - 1
The notation f(-9) simply tells you that you substitute -9 for x in the expression
f(x) = 2x-1
In this case the substitution of -9 for x leads to:
f(-9) = 2(-9) - 1 = -18 - 1 = -19
so f(-9) = -19